h=(4l-10)/5n
Solving for L

Answer by NerdvanaGirl(8)

(5hn+10)/4=L
Here are the steps:
h=(4L-10)/5n
Multiply both sides by 5n
h*5n = 4L-10 (because the 5n cancels)
Rewrite/simplify
5hn=4L-10
Add 10 to both sides
5hn+10=4L
Divide by 4
(5hn+10)/4=L